Suppose 𝑓 is odd and differentiable everywhere. Prove that for every positive



number 𝑏, there exists a number 𝑐 in (βˆ’π‘, 𝑏) such that 𝑓 β€²(𝑐) = 𝑓(𝑏)/𝑏.

1
Expert's answer
2022-03-31T02:51:58-0400

"Since\\,\\,f\\,\\,is\\,\\,odd, f\\left( -b \\right) =-f\\left( b \\right) \\\\By\\,\\,Lagrange\u0091s\\,\\,theorem\\\\\\frac{f\\left( b \\right) -f\\left( -b \\right)}{b-\\left( -b \\right)}=f'\\left( c \\right) ,c\\in \\left( -b,b \\right) \\\\\\frac{2f\\left( b \\right)}{2b}=f'\\left( c \\right) \\Rightarrow f'\\left( c \\right) =\\frac{f\\left( b \\right)}{b}"